OpenStudy (anonymous):

n a regular deck of 52 cards, there is a .25 probability of randomly selecting a card in the heart suit. Would the experimental probability be higher, lower, or equal to .25?

OpenStudy (anonymous):

i think it would be equal to right?

OpenStudy (turingtest):

that's what I would think too.

OpenStudy (zarkon):

it could be higher/lower or equal

OpenStudy (turingtest):

but it SHOULD be equal over enough trials.

OpenStudy (zarkon):

it should be close to .25 when a large number of trials is done

OpenStudy (unklerhaukus):

is the experiment legitimate /

OpenStudy (turingtest):

Are we in Vegas or Atlantic City?

OpenStudy (anonymous):

huh?

OpenStudy (unklerhaukus):

well if your holding two hearts and there is two on the board, and you just need that last one for the flush of the river, ...... your chances are certainly less than 1/4

OpenStudy (unklerhaukus):

for the flush on the river*

OpenStudy (anonymous):

true

OpenStudy (unklerhaukus):

in a valid experiment ,i.e. shuffled deck, enough samples, i cannot see why you wouldn't get the expectation value

OpenStudy (turingtest):

mine was just a bad joke...

OpenStudy (zarkon):

what if the theoretical probability is an irrational number. The experimental probability is always a rational number so it will only equal the theoretical probability in the limit

OpenStudy (unklerhaukus):

can you give example of a experiment involving a theoretical irrational probability

OpenStudy (turingtest):

Good question Unk I wanna know too

OpenStudy (zarkon):

Let \[X\] have a continuous uniform distribution. What is the probability thet X is less than 1/pi \[\mathbb{P}\left(X<\frac{1}{\pi}\right)\]

OpenStudy (turingtest):

yes but what sort of experiment would this formula correspond to?

OpenStudy (zarkon):

I should have specified the support...lets use uniform (0,1) randomly generate a value from a uniform (0,1). That is the experiment

OpenStudy (turingtest):

hmmm... is this a computer science type of experiment? I'm not familiar with this terminology "use uniform (0,1) randomly generate a value from a uniform (0,1)." This may just be over my head, I've never taken probability or done much computer science, but thank's for the lead, I'll look into it.

OpenStudy (unklerhaukus):

cards?

OpenStudy (zarkon):

with cards...no...but this statement "in a valid experiment ,i.e. shuffled deck, enough samples, i cannot see why you wouldn't get the expectation value " is still false

OpenStudy (unklerhaukus):

can you explain to me why "?

OpenStudy (zarkon):

it sounds like you are saying that if we repeated the experiment enough times we would be guaranteed to have the experimental prob=theoretical prob. is that what you are saying?

OpenStudy (unklerhaukus):

if we repeat a valid experiment,yes,

OpenStudy (zarkon):

how large would the number trials have to be?

OpenStudy (unklerhaukus):

that depends of the uncertainty in the measurement

OpenStudy (turingtest):

like you said, they should converge to the theoretical probability as the number of trials approaches infinity

OpenStudy (zarkon):

what about for this problem?

OpenStudy (zarkon):

that is true...they are equal in the limit, but for any finite number of trials we are not guaranteed to have equality

OpenStudy (turingtest):

well then in order to answer the original question we need to know how they defined "experimental probability" as opposed to "theoretical probability"

OpenStudy (unklerhaukus):

well if you get to the end of the 52 two cards i bet you'll have 1/4 hearts.... joking.

OpenStudy (jamesj):

This conversation is a good example of why we need a new system on openstudy, where people like Zarkon are known as 'people who know what the hell they're talking about' In practice, over a sample of n draws, the number of times a heart is drawn, h(n), could be more or less or equal to n/4. However is also is true is that \[\lim_{n \rightarrow \infty} \frac{h(n)}{n} = \frac{1}{4}\] But that doesn't mean that h(1) = 1/4 (in fact, it can't, as h(.) is integer valued); nor does it mean h(4) = 1 with absolute certainty. If you're really in any doubt about this. Take a deck of cards, shuffle them thoroughly, pick a card, note its suit, replace the card, reshuffle thoroughly and repeat a few times.

OpenStudy (zarkon):

I'm pretty sure JamesJ would fall into that category too.

OpenStudy (unklerhaukus):

experiments cannot have absolute certainty . it is a physical impossibility. going back to measuring experimentally a theoretical irrational numbers. i was hoping you were gonna say something like, Measuring the coastline of a roughly country. The Value determined would be dependent on the size of your measuring stick, as coastlines are not square-edged, a 10 km ruler or a 30 cm ruler would not agree on the coastline as all measuring sticks make approximations around the curves. With a small measuring stick a larger coastline is measured. The true coastline length is kinda like a fractal, so i guess the theoretical expectation value would be a function of the ruler

OpenStudy (zarkon):

Apparently you are hung up on actually performing the experiment...how about this...make a spinner...like one a child would use for a game. the spinner is broken up into 4 zones. It looks like each are the same size but we know that it is not possible to be that exact...therefore the theoretical probabilities will not be exactly .25 for each zone. there is a chance (I'd say with probability 1) that the regions have a probability that is an irrational number......